# Diffusion on language model encodings for protein sequence generation

### v2

@@ Abstract
Protein sequence design has seen significant advances through discrete diffusion and autoregressive approaches, yet the potential of continuous diffusion remains underexplored. Here, we present DiMA, a latent diffusion framework that operates on protein language model representations. Through systematic exploration of architectural choices and diffusion components, we develop a robust methodology that generalizes across multiple protein encoders ranging from 8M to 3B parameters. We demonstrate that our framework achieves consistently high performance across sequence-only (ESM-2, ESMc), dual-decodable (CHEAP), and multimodal (SaProt) representations using the same architecture and training approach. We extensively evaluate existing methods alongside DiMA using multiple metrics across two protein modalities, covering quality, diversity, novelty, and distribution matching of generated proteins. DiMA consistently produces novel, high-quality and diverse protein sequences and achieves strong results compared to baselines such as autoregressive, discrete diffusion and flow matching language models. The model demonstrates versatile functionality, supporting conditional generation tasks including protein family-generation, motif scaffolding and infilling, and fold-specific sequence design. This work provides a universal continuous diffusion framework for protein sequence generation, offering both architectural insights and practical applicability across various protein design scenarios.
Machine Learning, ICML
@@ 1 Introduction
In this study, we develop DiMA, a new latent diffusion model that operates on pLM representations. We demonstrate that continuous diffusion on protein embeddings enables effective sequence and structure generation across multiple tasks and encoder architectures. Our key contributions are as follows:
@@ 2 Continuous diffusion on LM representations of protein sequences
We utilize a pre-trained transformer-based pLM as an encoder with ESM-2 (Lin et al., 2023a) being the default choice unless otherwise specified. The encoder maps the sequence of discrete amino acids $y=[y_{1},...,y_{s}]$ of length $s$ to the latent vectors $x=[x_{1},...,x_{s}]\in\mathbb{R}^{s\times d}$, $x=\mathcal{E}(y)$. We employ dimension normalization to encourage each component of a single vector in the sequence $x$ to have zero mean and unit variance $z_{0}=Normalize(x)$. This transformation allows us to adapt the discrete protein input to a standard Gaussian diffusion framework.
[…]
@@
where $d$ is a hyperparameter reflecting the schedule’s rate. The larger the value of $d$, the greater the data corruption rate. We utilize a heuristic approach based on the observation that the reconstruction loss should exhibit an approximately linear increase over diffusion time (Figure 1).
@@
$\mathcal{L}(\theta)=\mathop{\mathbb{E}}_{\varepsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim U[0;1]}\left[||z_{0}-\hat{z}_{\theta}(z_{t},t,\text{SG}[\hat{z}_{0,t}])||^{2}\right]$
where $\text{SG}[\cdot]$ denotes the stop-gradient operation. Unlike (Chen et al., 2022), we apply a linear transformation to $\hat{z}_{0,t}$ and incorporate it into each transformer block’s input.
@@
Our architecture utilizes the ESM-2 decoder, which was pre-trained alongside the encoder on masked language modeling objectives. We found that additional fine-tuning of the decoder, specifically for amino acid reconstruction, improves sequence generation accuracy from latent representations $x$ during inference. The decoder maintains a simple architecture with a single linear layer.
@@ 3 Experiments
3.1 Evaluation Metrics
@@
3.2 Denoiser Component Analysis
In this part of our study, we utilize the ESM-8M encoder and DiMA-35M model for our experiments. To assess the contribution of the proposed design choices to the performance of DiMA, we train several models from scratch with the following modifications:
  • Removing the long skip-connections between the shallow and the deep transformer blocks.
  • Using time conditioning through admixing the time embeddings to the corrupted latent vectors of amino acids instead of employing a dedicated time layer before each transformer block.
  • Omitting the transformer encoder (ESM-2), retaining only its embedding matrix.
  • Training the model without self-conditioning.
  • Training models with linear and cosine noise schedule.
  • Training models with padding reconstruction and without prior length sampling.
  • Omitting finetuning the decoder.
  • Using flow matching paradigm in our latent generative model.
Table 1 demonstrates that each proposed feature contributes significantly to the model’s performance individually. The most substantial decrease in both the quality and distribution similarity of the generated sequences occurs in the ablated models without the ESM-2 encoder, padding omitting and length sampling, and when trained without self-conditioning. Removing skip-connections and time layers results in a less pronounced impact, but still significant decrease in repetitions of generated sequences and a slight improvement in overall quality.
To ablate the impact of the sd-$10$ noise schedule, we train our diffusion model with standard linear and cosine schedules, leaving other parameters intact. We find that sd-$10$ significantly outperforms the cosine schedule in both quality and distribution similarity. It also achieves less expressed but better results than the linear schedule.
Detailed results are provided in the Appendix E.1.
3.3 Comparison Across Generative Paradigms
@@ 3.3.1 Biological relevance
3.4 Representation Space Scaling
The AFDBv4-90 dataset represents a carefully curated subset of UniRef50. Comprising 2.2 million protein sequences with $>90\%$ AlphaFold2-predicted structural confidence, this dataset by design excludes intrinsically disordered proteins and low-entropy sequences, ensuring a high-quality protein corpus. As a consequence, this curation allows us to use pLDDT as a reliable, consistent measure of protein structural quality.
@@ 3.5 Comparison with Large Pretrained Models
We evaluate DiMA-35M, trained on ESM-3B encodings, against state-of-the-art pre-trained protein models RITA (Hesslow et al., 2022), ProtGPT2 (Ferruz et al., 2022), ProGen2 (Madani et al., 2023), EvoDiff (Alamdari et al., 2023), ProLLAMA (Lv et al., 2024), DPLM (Wang et al., 2024), Chroma (Ingraham et al., 2023), Multiflow (Campbell et al., 2024), RFDiffusion (Watson et al., 2023) across multiple configurations. For all models, we use the authors’ recommended sampling parameters to ensure a fair comparison. Only models with publicly available pre-trained weights are considered, ensuring transparency and reproducibility. Figure 2 illustrates the relationship between quality, diversity, and model size.
@@ 3.6.3 Fold-conditioned Generation
We finetune DiMA on the CATH S40 non-redundant dataset ( 27k proteins) and evaluate performance on a hold-out set of 100 structures. For each structure, we generate 10 proteins and assess their similarity to the target fold using the TM-score. Following the protocol of (Watson et al., 2023), we consider generation successful if at least one design achieves TM-score $>0.5$, indicating significant structural similarity. For comparison, we evaluate against RFDiffusion, a specialized structure generation model capable of fold-conditioning.
@@ 4 Conclusion
In this paper, we introduce DiMA, a continuous diffusion-based model for protein sequence generation that operates within the space of protein model representations. A comprehensive ablation study quantitatively verifies the impact of DiMA’s architectural features and design choices on its performance. Through extensive experiments, we evaluate the quality, diversity, distribution similarity, and biological relevance of the generated sequences. The results demonstrate that DiMA achieves comparable protein generation quality with multibillion models while utilizing a hundred times fewer parameters. Overall, this findings suggest that DiMA models are capable of generating diverse variants of natural-like proteins. The framework presented in this study provides a foundation for future research in protein generation.
5 Impact statement
@@ B.1 Quality
Here, $\mathcal{P}_{ESM-2}(S)$ represents the pseudoperplexity of sequence $S$, $|S|$ denotes the length of sequence $S$, $s_{i}$ is the $i$-th token in the sequence, $S_{\backslash i}$ represents the sequence without the $i$-th token, and $\Theta_{ESM-2}$ denotes the parameters of the ESM-2 model.
@@
[…]
@@
Distributional similarity FD-seq OT-seq MMD-seq FD-struct OT-struct MMD-struct
Quality ProGen-2 ppl ESM-2 pppl BLAST scPerplexity pLDDT TM-score scPerplexity
Diversity Rep CD
Novelty Novelty
@@ B.2 Diversity
Rep. Rep quantifies the internal diversity of generated sequences by assessing the prevalence of repeated subsequences, it is calculated as $\textbf{Rep}(y)=1-\prod_{n\in\{8,16,32,64\}}\frac{|\text{\# of unique n-subseqin }y|}{|\text{\# of n-subseq in }y|}$, where $y$ is a set of generated proteins. n-subseq means the subsequence of consecutive amino acids of length n.
@@
UMAP. To visually represent the distribution of generated sequences across PC, we trained UMAP on all sequences from PC for all models (parameters - n_neighbors - 25 and min_dict - 0.5). The UMAP plots in Figures  6 and  6 show that despite the fact that the diversity metrisc of the DIMA w/o self-conditioning are higher, DIMA with self-condition has the same coverage on the SwissProt (and even more coverage on AFDBv4-90). This and the fact that DIMA is closer to the dataset in terms of distribution learning metrics shows that DIMA w/o self condition achieved better diversity by generating sequences that greatly differ from those from the dataset.
@@ B.3 Distribution similarity
$d(X_{1},X_{2})^{2}=||\mu_{1}-\mu_{2}||^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2\sqrt{\Sigma_{1}\Sigma_{2}})$ (6)
@@ Appendix D Related Work
Diffusion generative models, introduced by Sohl-Dickstein et al. (2015), have gained attention for their remarkable results in image (Ho et al., 2020; Song et al., 2020b, a), and speech generation (Chen et al., 2020; Popov et al., 2021). Due to their impressive generative quality, some studies have extended the application of diffusion models to the text domain. Hoogeboom et al. (2021) and Austin et al. (2021) proposed multinomial diffusion for discrete data corruption. Subsequently, other works (Li et al., 2022; Lin et al., 2023b; Gulrajani & Hashimoto, 2023; Han et al., 2022; Strudel et al., 2022; Gao et al., 2022) adapted Gaussian diffusion to sequence learning by embedding discrete data into continuous space. Yuan et al. (2022) extended the text diffusion model to the sequence-to-sequence setting. Ye et al. (2023) conducted a study on the discrepancy of the text embedding space, demonstrating that the diffusion task at small noise scales is trivial. Zhang et al. (2023b) implemented latent text diffusion inside a VAE with an autoregressive decoder. Lovelace et al. (2022) utilized diffusion models to generate a fixed-length latent representation, mapped into a high-dimensional space with the reconstruction network before being fed into an autoregressive decoder to generate text.

### v4

@@ Abstract
Protein sequence design has seen significant advances through discrete diffusion and autoregressive approaches, yet the potential of continuous diffusion remains underexplored. Here, we present DiMA, a latent diffusion framework that operates on protein language model representations. Through systematic exploration of architectural choices and diffusion components, we develop a robust methodology that generalizes across multiple protein encoders ranging from 8M to 3B parameters. We demonstrate that our framework achieves consistently high performance across sequence-only (ESM-2, ESMc), dual-decodable (CHEAP), and multimodal (SaProt) representations using the same architecture and training approach. We extensively evaluate existing methods alongside DiMA using multiple metrics across two protein modalities, covering quality, diversity, novelty, and distribution matching of generated proteins. DiMA consistently produces novel, high-quality and diverse protein sequences and achieves strong results compared to baselines such as autoregressive, discrete diffusion and flow matching language models. The model demonstrates versatile functionality, supporting conditional generation tasks including protein family-generation, motif scaffolding and infilling, and fold-specific sequence design. This work provides a universal continuous diffusion framework for protein sequence generation, offering both architectural insights and practical applicability across various protein design scenarios. Code is released at GitHub.
Machine Learning, ICML, protein design, diffusion
Figure 1: DiMA. The framework consists of three main components: (1) a pre-trained protein language model encoder that maps amino acid sequences to continuous latent representations, (2) a diffusion denoiser that generates latent vectors from Gaussian noise, and (3) sequence and structure decoders that reconstruct amino acid sequences and protein structures from the generated latent representations. During training, the model learns to denoise corrupted protein representations. During inference, the framework supports both unconditional generation and conditional generation tasks including motif scaffolding, fold conditioning, and family-specific generation. The approach enables joint sequence-structure generation while operating entirely in continuous latent space.
@@ 1 Introduction
In this study, we develop DiMA, a new latent diffusion model that operates on pLM representations. We demonstrate that continuous diffusion on protein embeddings enables effective sequence and structure generation across multiple tasks and encoder architectures. DiMA addresses the limitations of previous continuous diffusion approaches that have been limited to specific representations by establishing a unified framework that generalizes across diverse protein encoders. By operating in the continuous latent space of these pre-trained encoders, our approach circumvents the challenges associated with discrete sequence modeling while maintaining the expressiveness needed for complex protein design tasks. The framework is designed to be encoder-agnostic, allowing it to benefit from advances in protein representation learning without requiring architectural modifications. Our key contributions are as follows:
@@ 2 Continuous diffusion on LM representations of protein sequences
Continuous vs. Discrete Diffusion for Proteins.
While discrete diffusion may appear more intuitive for sequence generation, continuous representations offer compelling advantages for protein modeling. Continuous encodings from protein language models capture rich semantic and structural information that has proven effective across diverse protein tasks, from representation learning to structure prediction. Recent studies have shown that continuous representations yield superior reconstruction accuracy compared to discrete alternatives (Lu et al., 2024), suggesting their potential for generative modeling.
Continuous diffusion offers several theoretical and practical advantages over discrete approaches such as direct application of established score-based techniques like classifier and classifier-free guidance without requiring discrete approximations. It also provides seamless integration with multimodal representations that jointly encode sequence and structure (e.g. CHEAP, SaProt). Also, continuous diffusion is more stable and efficient in training compared to discrete ones. However, the sequential and discrete nature of protein sequences presents unique challenges for continuous diffusion that necessitate careful adaptation of diffusion components to effectively capture the complexity of protein space.
@@
We utilize a pre-trained transformer-based pLM as an encoder with ESM-2 (Lin et al., 2023a) being the default choice unless otherwise specified. The encoder maps the sequence of discrete amino acids $y=[y_{1},...,y_{s}]$ of length $s$ to the latent vectors $x=[x_{1},...,x_{s}]\in\mathbb{R}^{s\times d}$, $x=\mathcal{E}(y)$. We apply normalization $z_{0}=Normalize(x)$ over the hidden dimension $d$: for each component $d_{i}$, we precompute the mean and variance over the training data and then apply normalization using these statistics to achieve zero mean and unit variance. This transformation allows us to adapt the discrete protein input to a standard Gaussian diffusion framework.
@@
where $d$ is a hyperparameter reflecting the schedule’s rate. The larger the value of $d$, the greater the data corruption rate. In this work we use $d=$ 10, hence the schedule is named tan-10. We utilize a heuristic approach based on the observation that the reconstruction loss should exhibit an approximately linear increase over diffusion time (Figure 2). The rationale behind the tan-10 schedule is discussed in Apeendix C.5.
[…]
@@
[…]
where $\text{SG}[\cdot]$ denotes the stop-gradient operation. Unlike (Chen et al., 2022), we apply a linear transformation to $\hat{z}_{0,t}$ and incorporate it into each transformer block’s input. The effect of self-conditioning rate on the quality-diversity trade-off is depicted on Figure 12
@@
Our architecture utilizes the ESM-2 decoder, which was pre-trained alongside the encoder on masked language modeling objectives. We found that additional fine-tuning of the decoder, specifically for amino acid reconstruction, improves sequence generation accuracy from latent representations $x$ during inference. The decoder maintains a simple architecture with a single linear layer. For CHEAP and SaProt representations, we fine-tune their corresponding pretrained decoders alongside the diffusion model, similar to our approach with ESM-2. We found that low-dimensional embeddings (e.g., ESM-2 8M, $d=$ 320) are less robust to small perturbations than higher-dimensional ones (ESM2/SaProt 650M, $d=$ 1280, CHEAP, $d=$ 1024). Fine-tuning the decoders helps minimize these effects during diffusion generation.
@@ 3 Experiments
3.1 Evaluation Metrics
3.2 Denoiser Component Analysis
In this part of our study, we utilize the ESM-8M encoder and DiMA-35M model for our experiments. To assess the contribution of the proposed design choices to the performance of DiMA, we train several models from scratch with the following modifications: removing the long skip-connections between the shallow and the deep transformer blocks; using time conditioning through admixing the time embeddings to the corrupted latent vectors of amino acids instead of employing a dedicated time layer before each transformer block; omitting the transformer encoder (ESM-2), retaining only its embedding matrix; training the model without self-conditioning; training models with linear and cosine noise schedule; training models with padding reconstruction and without prior length sampling; omitting finetuning the decoder; and using flow matching paradigm in our latent generative model.
@@ 3.2 Denoiser Component Analysis
Table 1 demonstrates that each proposed feature contributes significantly to the model’s performance individually. The most substantial decrease in both the quality and distribution similarity of the generated sequences occurs in the ablated models without the ESM-2 encoder, without length sampling, and when trained without self-conditioning. Removing skip-connections and time layers results in a less pronounced impact, but still significant decrease in repetitions of generated sequences and a slight improvement in overall quality.
To ablate the impact of the tan-$10$ noise schedule, we train our diffusion model with standard linear and cosine schedules, leaving other parameters intact. We find that tan-$10$ significantly outperforms the cosine schedule in both quality and distribution similarity. It also achieves less expressed but better results than the linear schedule. The detailed results are provided in the Appendix E.1.
3.3 Comparison Across Generative Paradigms
@@ 3.3.1 Biological relevance
3.4 Representation Space Scaling
The AFDBv4-90 dataset represents a carefully curated subset of UniRef50. Comprising 2.2 million protein sequences with $90\%$ AlphaFold2-predicted structural confidence, this dataset by design excludes intrinsically disordered proteins and low-entropy sequences, ensuring a high-quality protein corpus. As a consequence, this curation allows us to use pLDDT as a reliable, consistent measure of protein structural quality.
@@ 3.5 Comparison with Large Pretrained Models
We evaluate DiMA-35M, trained on ESM-3B encodings, against state-of-the-art pre-trained protein models, namely, RITA (Hesslow et al., 2022), ProtGPT2 (Ferruz et al., 2022), ProGen2 (Madani et al., 2023), EvoDiff (Alamdari et al., 2023), ProLLAMA (Lv et al., 2024), DPLM (Wang et al., 2024), Chroma (Ingraham et al., 2023), Multiflow (Campbell et al., 2024), RFDiffusion (Watson et al., 2023), and PLAID-100M (PLAID). For all models, we use the authors’ recommended sampling parameters to ensure a fair comparison. However, for autoregressive models ProGen2 and ProLLAMA, which show suboptimal quality and collapse to highly repetitive sequences on default settings, we performed grid searches to identify optimal temperature and top-p values. We consider only models with publicly available pre-trained weights, ensuring transparency and reproducibility. Figure 3 illustrates the relationship between quality, diversity, and model size.
@@ 3.6.3 Fold-conditioned Generation
We finetune DiMA on the CATH S40 non-redundant dataset ( 27k proteins) and evaluate performance on a hold-out set of 100 structures. For each structure, we generate 10 proteins and assess their similarity to the target fold using the TM-score. Following the protocol of (Watson et al., 2023), we consider generation successful if at least one design achieves TM-score >0.5>0.5, indicating significant structural similarity. For comparison, we evaluate against RFDiffusion, a specialized structure generation model capable of fold-conditioning.
@@ 4 Conclusion
In this paper, we introduce DiMA, a continuous latent diffusion framework for protein sequence generation that operates effectively across diverse protein language model representations. Through systematic exploration of architectural choices and diffusion components, we establish a robust methodology that generalizes from sequence-only encoders to multimodal representations spanning 8M to 3B parameters.
DiMA demonstrates versatile functionality across multiple protein design scenarios, including unconditional generation, family-specific design, motif scaffolding, sequence infilling, and fold-conditioned generation. The framework achieves competitive performance with significantly larger models while maintaining computational efficiency, as the encoder becomes dispensable during inference.
This work provides both architectural insights and practical applicability for protein sequence generation, offering a unified approach that bridges sequence and structure generation through continuous diffusion. The systematic methodology we establish creates a foundation for future developments in computational protein design, demonstrating how domain-specific adaptations can unlock the potential of diffusion models for biological sequence generation.
Impact statement
@@ B.1 Quality
[…]
@@
$\text{TM-score}=\frac{1}{L_{\text{target}}}\sum_{i=1}^{L_{\text{query}}}\frac{1}{1+\left(\frac{d_{i}}{d_{0}(L_{\text{target}})}\right)^{2}}$ (5)
@@
Distributional
similarity
FD-seq
OT-seq
MMD-seq
FD-struct
OT-struct
MMD-struct
Quality
ProGen-2 ppl
ESM-2 pppl
BLAST
scPerplexity
pLDDT
TM-score
scPerplexity
Diversity
Rep
CD
Novelty
Novelty
@@ B.2 Diversity
[…]
@@
UMAP. To visually represent the distribution of generated sequences across PC, we trained UMAP on all sequences from PC for all models (parameters - n_neighbors - 25 and min_dict - 0.5). The UMAP plots in Figures  7 and  7 show that despite the fact that the diversity metrisc of the DIMA w/o self-conditioning are higher, DIMA with self-condition has the same coverage on the SwissProt (and even more coverage on AFDBv4-90). This and the fact that DIMA is closer to the dataset in terms of distribution learning metrics shows that DIMA w/o self condition achieved better diversity by generating sequences that greatly differ from those from the dataset.
@@ B.3 Distribution similarity
[…]
@@ C.3 Length Determination
C.4 Parameters for Navigating the Quality-Diversity Trade-off
Figure 11: Effect of the number of sampling steps on DiMA-35M generation quality. Increasing the number of diffusion steps from 250 to 2000 shows a trade-off between quality and diversity. Top: structural quality (pLDDT) improves with more sampling steps, approaching natural protein level (gray dashed line). Middle: sequence diversity ($CD_{0.5}$) slightly decreases with more steps but remains consistently high (>>0.97). Bottom: novelty decreases with increased sampling steps, indicating convergence toward more conservative protein designs. Each point represents evaluation on 2048 protein sequences generated with the specified number of steps.
Figure 12: Effect of self-conditioning rate parameter ($w$) on DiMA-35M performance metrics. In self-conditioned generation $x_{0}=\hat{x}_{0}+w*(x_{0}-\hat{x}_{0})$, the parameter $w$ controls how strongly the model relies on its own previous predictions during the diffusion process. Top: structural quality (pLDDT) increases with higher $w$ values, approaching natural protein levels (gray dashed line) at $w\geq$ 0.9. Middle: sequence diversity ($CD_{0.5}$) decreases as $w$ increases, with noticeable reduction at $w>$ 1.0. Bottom: novelty follows a similar trend, showing higher values at lower $w$. Each point represents evaluation on 2048 protein sequences generated with 2000 sampling steps and the specified $w$ value.
C.5 The Rationale Behind the tan-10 Schedule
The key issue with standard noise schedules (linear, cosine) is that they corrupt data very gradually at small timesteps (Figure 2). This means the model spends significant training resources on nearly trivial denoising tasks where input and target are almost identical. Our approach ensures the difficulty of the denoising task increases steadily with each timestep. By designing a noise schedule where reconstruction loss grows linearly with diffusion time, we create learning problems of incrementally increasing difficulty, allowing the model to make consistent progress throughout training rather than facing an abrupt jump in task complexity.
[…]
Figure 13: Structural stability of protein representations under noise. Root Mean Square Deviation (RMSD) between original protein structure and structure reconstructed from noisy CHEAP representations at different diffusion timesteps.
Figure 14: Sequence reconstruction accuracy under noise. Percentage of amino acids correctly reconstructed from noisy CHEAP representations at different diffusion timesteps.
@@ Appendix D Related Work
Diffusion generative models, introduced by Sohl-Dickstein et al. (2015), have gained attention for their remarkable results in image (Ho et al., 2020; Song et al., 2020b, a), and speech generation (Chen et al., 2020; Popov et al., 2021). Due to their impressive generative quality, some studies have extended the application of diffusion models to the text domain. Hoogeboom et al. (2021) and Austin et al. (2021) proposed multinomial diffusion for discrete data corruption. Subsequently, other works (Li et al., 2022; Lin et al., 2023b; Gulrajani & Hashimoto, 2023; Han et al., 2022; Strudel et al., 2022; Gao et al., 2022; shabalin2025tencdm) adapted Gaussian diffusion to sequence learning by embedding discrete data into continuous space. Yuan et al. (2022) extended the text diffusion model to the sequence-to-sequence setting. Ye et al. (2023) conducted a study on the discrepancy of the text embedding space, demonstrating that the diffusion task at small noise scales is trivial. Zhang et al. (2023b) implemented latent text diffusion inside a VAE with an autoregressive decoder. Lovelace et al. (2022) utilized diffusion models to generate a fixed-length latent representation, mapped into a high-dimensional space with the reconstruction network before being fed into an autoregressive decoder to generate text.